OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x) @jigglypuff314

jigglypuff314 (jigglypuff314):

alright :) so inverse means that you switch the x for a y and a y for the x so x = 7y - 13 solve for y :)

OpenStudy (anonymous):

+13 to both sides?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

13x = 7y?

OpenStudy (anonymous):

are you looking for the inverse or -1?

OpenStudy (anonymous):

im not sure this is what the problem says

OpenStudy (anonymous):

funny, i did not see an option A

OpenStudy (anonymous):

y= 7x-13 y+13 = 7x x = (y+13)/7 now to find inverse function just replace x and y y= (x+13)/7 so your inverse function is f−1(x)=x+137 got this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how would you solve this? Let f(x) = x2 - 3x - 7. Find f(-3). would it be -3 = -3(2) - -3(-3)?

jigglypuff314 (jigglypuff314):

mmm f(-3) means x = -3 so (-3)^2 - 3(-3) - 7

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

9 - -9 - 7 = 11?

OpenStudy (anonymous):

I actually just wrote this on someone else's post, and copy pasted here, since it fits the question, anyways sorry to intrude jiggly, continue.

jigglypuff314 (jigglypuff314):

lol it's fine iambatman xP I'm kinda lost... but yeah 11 turtle

OpenStudy (the_fizicx99):

How're you lost :o

OpenStudy (anonymous):

:) good job turtle

OpenStudy (anonymous):

thank you:) what about Let f(x) = 2x - 6. Solve f-1(x) when x = 2 would i do 2 = 2(2) - 6?

jigglypuff314 (jigglypuff314):

first get f-1(x) so x = 2y - 6 solve for y

OpenStudy (anonymous):

+ 6 on both sides?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

6x = 2y do i divide 6 by 2?

jigglypuff314 (jigglypuff314):

mmmm x + 6 = 2y divide both sides by 2

OpenStudy (anonymous):

would it be 3?

jigglypuff314 (jigglypuff314):

not quite... mmm they want you to find x= 2 right? so 2+6 = 2y solve for y ?

OpenStudy (anonymous):

oh ok so would it be 4?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

ok thank you last question for the night i promise Let f(x) = 2x + 2. Solve f-1(x) when x = 4.

jigglypuff314 (jigglypuff314):

we've done a couple of f-1(x) type problems now, can you try setting it up yourself this time? :)

OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4 -1(4) =2(4) +2?

OpenStudy (anonymous):

-4 = 8 +2 add 2 to both sides -2 = 8 now divide -4?

jigglypuff314 (jigglypuff314):

I'm confused with where you're going with this :/ x = 2y + 2 when you flip the x and y then x = 4 so 4 = 2y + 2 solve for y

OpenStudy (anonymous):

would it be 2?

jigglypuff314 (jigglypuff314):

mmm not quite :)

OpenStudy (anonymous):

thank you so much for your help!!! i have to go to dinner thank you:)

jigglypuff314 (jigglypuff314):

I'm glad I could help :)